OpenStudy (anonymous):

log[2](x)((1)/(9))=-2 answer plz the viarable is x

OpenStudy (marissalovescats):

Could you write this out to look more simple? :P

hartnn (hartnn):

\(\large \log_{2} \dfrac{1}{9}=-2\) where is the variable ?? O.o

Directrix (directrix):

Is there a variable missing: log[2]((1)/(9))=-2 answer plz

OpenStudy (jdoe0001):

$$ \huge log_2\pmatrix{\cfrac{1}{9}}= -2 $$

Directrix (directrix):

That yields 1/4 = 1/9, doesn't it?

OpenStudy (jhannybean):

They don't equal,hence the missing variable.

hartnn (hartnn):

@sidney38

OpenStudy (marissalovescats):

I think this is like a rearrange kinda thing... So no variables in this log.

OpenStudy (jhannybean):

There needs to be because 1/4= 1/9 is a false statement.

OpenStudy (jhannybean):

The only way to make a true statement would be to solve for x.

ganeshie8 (ganeshie8):

variable is x yeah so varible is missing

OpenStudy (jdoe0001):

@sidney can you take a screenshot of it?

hartnn (hartnn):

^ you tagged different sidney :P

OpenStudy (jhannybean):

hahaha.

OpenStudy (jhannybean):

@sidney38

OpenStudy (jdoe0001):

ohh shoot, wrong sidney darn

hartnn (hartnn):

she messaged me that the variable is x maybe she is unable to post here....

OpenStudy (jhannybean):

refresh page...

OpenStudy (campbell_st):

it works if it a base 3 log.... otherwise its false.

OpenStudy (campbell_st):

\[\log_{3} \frac{1}{9} = \log_{3} 3^{-2} = -2\]

OpenStudy (jhannybean):

True enough.

hartnn (hartnn):

now i can see a 'x'

OpenStudy (jhannybean):

yes, the "x" would be the \(3^{-2}\)

ganeshie8 (ganeshie8):

\(\large \log_2{\frac{x}{9}}=-2 \)

OpenStudy (jhannybean):

\[\large 2^{-2} = \cfrac{x}{9}\]\[\large \cfrac{1}{4}=\cfrac{x}{9} \]

hartnn (hartnn):

\(\huge \log_ab=c \implies b=a^c\)

hartnn (hartnn):

\(\large \log_2{\frac{x}{9}}=-2\) \(\implies\) what jhannybean posted...

OpenStudy (jdoe0001):

$$ log_2\pmatrix{\cfrac{x}{9}}= -2\\ 2^{log_2\pmatrix{\cfrac{x}{9}}} = 2^{-2}\\ \text{using the log cancellation rule}\\ \color{blue}{a^{log_ax} = x}\\ 2^{-2} = \cfrac{x}{9}\\ \cfrac{1}{4}=\cfrac{x}{9} $$ as shown above by Jhannybean